A solution with the concentration of 1.53 mg/dL must be diluted to a concentration of 8.33 ug/dL . if the starting solution contains 5.5L, what is the final volume of the solution in liters after it reaches the desired 8.33ug/dL concentration?
1
Expert's answer
2017-12-11T07:38:11-0500
Solution
Find the mass of the compound in the solution:
m = 1.53 × 5.5 = 8.415 (mg) or 8 415 μg
If 8.33 μg must be in 1L,
Then 8 415 μg must be in = 1 010.2 L
Answer
1 010.2 L – the final volume of the solution in liters after it reaches the desired 8.33 μg/dL
concentration.